A company gives an aptitude test to potential employees. The scores are normally distributed with mean 250 and S.D 50.What is the probability that a score is selected at random will be Between 230 and 260.


1
Expert's answer
2021-05-30T23:53:21-0400

Let "X=" the score: "X\sim N(\mu, \sigma^2)"

Given "\mu=250, \sigma=50"


"P(230<X<260)=P(X<260)-P(X\leq 230)"

"=P(Z<\dfrac{260-250}{50})-P(Z\leq \dfrac{230-250}{50})"

"=P(Z<0.2)-P(Z\leq -0.4)"

"\approx0.5792597-0.34457823\approx0.234681"